Grade 11Magical Mathematics[Interesting Approach]

A trader buys goods at a 19% discount on the label price. If he wants to make a profit of 20% after allowing a discount of 10%, by what % should his marked price be greater than the original label price?

Shobha Sexana
17 Years agoGrade 11

1 Answer

ritu goyal
17 Years ago

Let the label price be = Rs.100. The trader buys at a discount of 19%.
Hence, his cost = 100 - 19 = 81.

He wants to make a profit of 20%. Hence his selling price = 1.2 (81) = 97.2

However, he wants to get this Rs.97.2 after providing for a discount of 10%. i.e. he will be selling at 90% of his marked price.
Hence, his marked price M = = 108 which is 8% more than the original label price.